OpenStudy (darkigloo):

How do I find empirical formula? Hydroquinone acid C=65.4g H=5.5g O=29.1g

OpenStudy (cuanchi):

divide the mass of each element by the molecular weight to find the moles of each.

OpenStudy (darkigloo):

Here is what I did but it is wrong. The answer is supposed to be C3H402. C=65.4/12 = 5.45 / 1.82 = 3 H= 5.5/1 = 5.5 / 1.82 = 3 0 = 29.1 / 16 = 1.82 / 1.82 = 1 So i got C3H3O as my answer. What did i do wrong?

OpenStudy (cuanchi):

then divide the number moles by the smallest one, in this case is going to be the number of moles of Oxygen

OpenStudy (cuanchi):

yes I think you are right, I got the same values

OpenStudy (cuanchi):

who said is C3H4O2? http://en.wikipedia.org/wiki/Hydroquinone molecular formula is C6H6O2

OpenStudy (darkigloo):

:O wow so i am right?! thank you!

OpenStudy (darkigloo):

There's another question that I got wrong for sure. Can I show you?

OpenStudy (cuanchi):

sure

OpenStudy (darkigloo):

acrylic acid C=50g /12 = 4.17 / 2.78 = 2 H= 5.6g /1 = 5.6 / 2.78 = 2 O= 44.4g / 16 = 2.78 / 2.78 = 1 I got C2H20 but the answer should be C3H4O2

OpenStudy (cuanchi):

yea that is wrong, you have to consider the C 4.17/2.78 =1.5 so then multiply everything times 2 to make whole numbers C3H4O2

OpenStudy (darkigloo):

Oh I see now. Thank you for your help. :)

OpenStudy (cuanchi):

np